OpenStudy (anonymous):

find the solution to the system of equations by using either graphing or substitution. y = 6 - x and y = x -2 a. (2,4) b. (-4,2) c. (4,2) d. no solutions

OpenStudy (anonymous):

so put 1 of the equations as y

OpenStudy (anonymous):

or like this 6-z=z-2

OpenStudy (anonymous):

then solve / simplify

OpenStudy (anonymous):

ok so it would be the first one?

OpenStudy (anonymous):

maybe

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y = 6 - x }\) \(\large\color{black}{ \displaystyle y = x -2 }\) the first equation tells you that y is the same thing as 6-x, so you can replace the y in the second equation by 6-x. your first equation \(\large\color{red}{ \displaystyle y = 6 - x }\) your second equation \(\large\color{black}{ \displaystyle y = x -2 }\) replace the y with 6-x in the second equation \(\large\color{black}{ \displaystyle 6-x = x -2 }\) (this is a substitution)

OpenStudy (anonymous):

yep yep

OpenStudy (anonymous):

thanks @SolomonZelman

OpenStudy (anonymous):

<3

OpenStudy (solomonzelman):

do you get the substitution method in this case, @120hoho, or not ?

OpenStudy (anonymous):

he said he thinks its the 1st one

OpenStudy (solomonzelman):

i doubt that the first answer is correct..... lets solve the problem and lets know for sure, okay?

OpenStudy (anonymous):

@SolomonZelman u got this right cuz i gtg

OpenStudy (solomonzelman):

Hoho, did you understand the substitution method that I explained?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

ok, so at this point we have: \(\large\color{black}{ \displaystyle 6-x = x -2 }\)

OpenStudy (solomonzelman):

you need to {solve for/ isolate the} x.

OpenStudy (solomonzelman):

lets add x to both sides. your equation right now: \(\large\color{black}{ \displaystyle 6-x = x -2 }\) adding x to both sides (as shown in blue) \(\large\color{black}{ \displaystyle 6-x \color{blue}{+x} = x -2\color{blue}{+x} }\)

OpenStudy (solomonzelman):

then tell me what is going to happen (what you get) on each side

OpenStudy (anonymous):

double x's

OpenStudy (solomonzelman):

more precisely please. What do you get on the right side? What do you get on the left side?